# Do We Need All the Synthetic Data? Towards Targeted Synthetic Image Augmentation via Diffusion Models

### v1

@@ A.1 Full Gradient Formulas & Useful Lemmas
=$\overset{(ii)}{=}\frac{3}{N}l_{i}^{(t)}\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle^{2}\left(1-\frac{3\rho^{(t)}}{N}l_{i}^{(t)}\|{\bm{\xi}}_{i}\|^{2}|\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle|\right)^{2}\|\bm{\xi}_{i}\|^{2},$ (9)
where (i) follows from Eq. 4 and (ii) is similar to [30], which claims that we can use the approximation for the logits in early training ($l_{i}^{(t)}=l_{i,\bm{\epsilon}}^{(t)}$, see proof of their Lemma A.17). Now with a sufficiently small $\rho$ that satisfies Eq. 3, the bound in 6 yields
$|\langle\nabla_{{\bm{w}}_{j,\bm{\epsilon}}}{\mathcal{L}}({\bm{W}}^{(t)}\!+\bm{\epsilon}^{(t)}),\bm{\xi}_{i}\rangle|$ $=\frac{3}{N}l_{i}^{(t)}{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle^{2}\underbrace{\left(1-\frac{3\rho^{(t)}}{N}l_{i}^{(t)}\|{\bm{\xi}}_{i}\|^{2}|\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle|\right)^{2}}_{\text{$\in(0,1)$}}\|\bm{\xi}_{i}\|^{2}$
$<\frac{3}{N}l_{i}^{(t)}{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle^{2}\|\bm{\xi}_{i}\|^{2}=|\langle\nabla_{{\bm{w}}_{j}}{\mathcal{L}}({\bm{W}}^{(t)}),\bm{\xi}_{i}\rangle|.$
@@
The special case of the theorem can be proved by directly setting the early training weights to be the initialization ${\bm{W}}^{(0)}$. Since ${\bm{W}}^{(0)}\sim\mathcal{N}(0,\sigma_{0}^{2})$ and $\bm{\xi}_{i}\sim\mathcal{N}(0,\sigma_{p}^{2}/d)$, with a large $N$, ${\mathcal{I}}_{j,+}^{(0)}$ and ${\mathcal{I}}_{j,-}^{(0)}$ will each contain roughly half of the noises, as $\text{sgn}(\langle{\bm{w}}_{j,\bm{\epsilon}}^{(0)},\bm{\xi}_{i}\rangle)=\text{sgn}(y_{i})$ has probability $0.5$ in this case.
@@
(Static noise sets) In our setting, with a sufficiently small learning rate $\rho$, for all $j\in[J]$ and all the early iterations $t=0,\dots,T$, we have that:
@@
This directly follows from the previous proofs, where starting from the intializations, certain noise alignments move towards 00 and others move away from 00 early in training. Hence, in this process, the signs do not change, and these sets always contain the same noises as $t$ progresses ∎
We now focus on the informal special case and show that in expectation, our requirement on the generation noise prevents $\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{G,(t)},{\bm{w}}_{j}^{(t)})$ from getting too large. In particular, we assume that the number of neurons $J$ is sufficiently large such that we can replace the noisy sigmoid logits with the noiseless version in Eq. 3:
$l_{i}^{(t)}=\text{sigmoid}\left(\sum_{j=1}^{J}-\beta_{d}^{3}\langle{\bm{w}}_{j}^{(t)},{\bm{v}}_{d}\rangle^{3}-y_{i}\langle{\bm{w}}_{j}^{(t)},\bm{\phi}_{i}\rangle^{3}\right)\approx l_{d}^{(t)}:=\text{sigmoid}\left(\sum_{j=1}^{J}-\beta_{d}^{3}\langle{\bm{w}}_{j}^{(t)},{\bm{v}}_{d}\rangle^{3}\right)$
or
$l_{i}^{(t)}=\text{sigmoid}\left(\sum_{j=1}^{J}-\beta_{e}^{3}\langle{\bm{w}}_{j}^{(t)},{\bm{v}}_{e}\rangle^{3}-y_{i}\langle{\bm{w}}_{j}^{(t)},\bm{\phi}_{i}\rangle^{3}\right)\approx l_{d}^{(t)}:=\text{sigmoid}\left(\sum_{j=1}^{J}-\beta_{d}^{3}\langle{\bm{w}}_{j}^{(t)},{\bm{v}}_{e}\rangle^{3}\right)$
depending on the feature. That is, the model has sufficient neurons $J$ to average out the second term in summation.
[…]
By our random initializations, we should again have $|{\mathcal{I}}_{j,+}^{G,(0)}|\approx|{\mathcal{I}}_{j,-}^{G,(0)}|$ and $|{\mathcal{I}}_{j,+}^{U,(0)}|\approx|{\mathcal{I}}_{j,-}^{U,(0)}|$, each randomly occupying around half of the noises. Lemma C.1 implies they continue to hold across early iterations. We now start with computing $\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{G,(0)},{\bm{w}}_{j}^{(0)})$: for simplicity, we let $\nabla_{j}(t):=\nabla_{{\bm{w}}_{j}^{(t)}}{\mathcal{L}}({\bm{W}}^{(t)})$,
$\ \ \ \ \ \ \frac{1}{|{\mathcal{I}}_{j,+}^{G,(0)}|}\left(\sum_{\begin{subarray}{c}i=1,\\ \bm{\xi}_{i}\in{\mathcal{I}}_{j,+}^{G,(0)}\end{subarray}}^{N}|\langle\nabla_{j}(0),{\bm{\xi}}_{i}\rangle|+\sum_{\begin{subarray}{c}i=N+1,\\ \bm{\gamma}_{i}\in{\mathcal{I}}_{j,+}^{G,(0)}\end{subarray}}^{N_{\text{new}}}|\langle\nabla_{j}(0),\bm{\gamma}_{i}\rangle|\right)$
[…]
${=}\ \frac{1}{|{\mathcal{I}}_{j,+}^{U,(0)}|}\left(\sum_{\begin{subarray}{c}i=1,\\ \bm{\xi}_{i}\in{\mathcal{I}}_{j,+}^{U,(0)}\end{subarray}}^{\alpha N}|\langle\nabla_{j}(0),{\bm{\xi}}_{i}\rangle|+\sum_{\begin{subarray}{c}i=\alpha N+1,\\ \bm{\xi}_{i}\in{\mathcal{I}}_{j,+}^{U,(0)}\end{subarray}}^{N}|\langle\nabla_{j}(0),{\bm{\xi}}_{i}\rangle|+\sum_{\begin{subarray}{c}i=N+1,\\ \bm{\gamma}_{i}\in{\mathcal{I}}_{j,+}^{G,(0)}\end{subarray}}^{N_{\text{new}}}|\langle\nabla_{j}(0),\bm{\gamma}_{i}\rangle|\right)\quad(ii)$
@@
$=\frac{1}{|{\mathcal{I}}_{j,+}^{U,(0)}|}\left(\sum_{\begin{subarray}{c}i=1,\\ \bm{\xi}_{i}\in{\mathcal{I}}_{j,+}^{U,(0)}\end{subarray}}^{\alpha N}|\langle\nabla_{j}(0),{\bm{\xi}}_{i}\rangle|+k\sum_{\begin{subarray}{c}i=\alpha N+1,\\ \bm{\xi}_{i}\in{\mathcal{I}}_{j,+}^{U,(0)}\end{subarray}}^{N}k|\langle\nabla_{j}(0),{\bm{\xi}}_{i}\rangle|\right)=\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{U,(0)},{\bm{w}}_{j}^{(0)}),$
@@
Then at a high level, the higher NoiseAlign at initialization for upsampling lets it start off with faster noise learning. This gets propagated to later iterations and in expectation,
$\langle{\bm{w}}_{j}^{G,(1)},\bm{\gamma}\rangle^{2}\|\bm{\gamma}\|^{2}<(k+1)\langle{\bm{w}}_{j}^{U,(1)},\bm{\xi}\rangle^{2}\|\bm{\xi}\|^{2}$
for the first iteration, where ${\bm{w}}_{j}^{G,(1)}$, ${\bm{w}}_{j}^{U,(1)}$ are the updated weights for the corresponding dataset. Given the higher NoiseAlign, this would hold as long as the higher gradient alignment updates the noise more for RHS upsampling. Subsequently, since Lemma C.1 implies that the noise multi-sets remain unchanged since initialization, the theorem statement follows inductively using the same bounding technique as above.
@@
$B_{1}=B\frac{\alpha N}{\alpha N+k(1-\alpha)N}=B\frac{\alpha}{\alpha+k(1-\alpha)}$ (11)
$B_{2}=B\frac{k(1-\alpha)N}{\alpha N+k(1-\alpha)N}=B\frac{k(1-\alpha)}{\alpha+k(1-\alpha)}$ (12)
@@
$=\frac{\sigma^{2}}{B}\left(1+\frac{k(k-1)(1-\alpha)}{(\alpha+k(1-\alpha))^{2}}\frac{B}{N}\right)=I$
@@ Appendix E Pseudocode
Algorithm 1 illustrates our method.
@@ F.1 Additional experimental settings
Datasets. The CIFAR10 dataset consists of 60,000 32 × 32 color images in 10 classes, with 6000 images per class. The CIFAR100 dataset is just like the CIFAR10, except it has 100 classes containing 600 images each. For both of these datasets, the training set has 50,000 images (5,000 per class for CIFAR10 and 500 per class for CIFAR100) with the test set having 10,000 images. Tiny-ImageNet comprises 100,000 images distributed across 200 classes of ImageNet [12], with each class containing 500 images. These images have been resized to 64×64 dimensions and are in color. The dataset consists of 500 training images, 50 validation images, and 50 test images per class.
Training on different datasets. From the setting in [2], we trained Pre-Activation ResNet18 on all datasets for 200 epochs with a batch size of 128. We used SGD with the momentum parameter 0.9 and set weight decay to 0.0005. We also fixed $\rho=0.1$ for SAM in all experiments unless explicitly stated. We used a linear learning rate schedule starting at 0.1. The learning rate is decayed by a factor of 10 after 50% and 75% epochs, i.e., we set the learning rate to 0.01 after 100 epochs and to 0.001 after 150 epochs.
Training with different architectures. We used the same training procedures for Pre-Activation ResNet18, VGG19, and DenseNet121. We directly used the official Pytorch [33] implementation for VGG19 and DenseNet121. For ViT [43], we adopted a Pytorch implementation at https://github.com/lucidrains/vit-pytorch. In particular, the hidden size, the depth, the number of attention heads, and the MLP size are set to 768, 8, 8, and 2304, respectively. We adjusted the patch size to 4 to fit the resolution of CIFAR10 and set both the initial learning rate and $\rho$ to 0.01.
Hyperparameters. For Useful [30], we adopt their default hyper-parameters for separating epochs and set the upsampling factor $k$ to 2 as it yields the best performance. For our method, the upsampling factor $k$ is set to $5$ for CIFAR10 and CIFAR100 while that of TinyImageNet is set to $4$. Table 1 summarizes the upsampling factor along with the denoising steps at which the synthetic images are used to augment the base training set.
@@
Generating synthetic data with diffusion model. We use an open-source text-to-image diffusion model, GLIDE [31] as our baseline for image generation. We directly used the official Pytorch implementation at https://github.com/openai/glide-text2im. While generating, we set the time steps as 100 and the guidance scale as 3.0. For a k-way classification, we input the class names $C=\{c_{1},\dots,c_{k}\}$ with prompt $l=^{\prime}\text{a photo of}\{c_{i}\}^{\prime}$.

### v2

@@ A.1 Full Gradient Formulas & Useful Lemmas
$\overset{(ii)}{=}\frac{3}{N}l_{i}^{(t)}\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle^{2}\left(1-\frac{3\rho^{(t)}}{N}l_{i}^{(t)}\|{\bm{\xi}}_{i}\|^{2}|\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle|\right)^{2}\|\bm{\xi}_{i}\|^{2},$ (9)
where (i) follows from Eq. 4 and (ii) uses the approximation for the logits in early training ($l_{i}^{(t)}=l_{i,\bm{\epsilon}}^{(t)}$). Now with a sufficiently small $\rho$ that satisfies Equation 3, the bound in 6 yields
$|\langle\nabla_{{\bm{w}}_{j,\bm{\epsilon}}}{\mathcal{L}}({\bm{W}}^{(t)}\!+\bm{\epsilon}^{(t)}),\bm{\xi}_{i}\rangle|$ $=\frac{3}{N}l_{i}^{(t)}\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle^{2}\underbrace{\left(1-\frac{3\rho^{(t)}}{N}l_{i}^{(t)}\|{\bm{\xi}}_{i}\|^{2}|\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle|\right)^{2}}_{\text{$\in(0,1)$}}\|\bm{\xi}_{i}\|^{2}$
$<\frac{3}{N}l_{i}^{(t)}\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle^{2}\|\bm{\xi}_{i}\|^{2}=|\langle\nabla_{{\bm{w}}_{j}}{\mathcal{L}}({\bm{W}}^{(t)}),\bm{\xi}_{i}\rangle|.$
@@
The special case of the theorem can be proved by directly setting the early training weights to be the initialization ${\bm{W}}^{(0)}$. Since ${\bm{W}}^{(0)}\sim\mathcal{N}(0,\sigma_{0}^{2})$ and $\bm{\xi}_{i}\sim\mathcal{N}(0,\sigma_{p}^{2}/d)$, with a large $N$, ${\mathcal{I}}_{j,+}^{(0)}$ and ${\mathcal{I}}_{j,-}^{(0)}$ will each contain roughly half of the noises, as $\text{sgn}(\langle{\bm{w}}_{j}^{(0)},\bm{\xi}_{i}\rangle)=\text{sgn}(y_{i})$ has probability $0.5$ in this case.
@@
(Static noise sets) In our setting, with a sufficiently small learning rate $\eta$, for all $j\in[J]$ and all the early iterations $t=0,\dots,T$, we have that:
@@
This directly follows from the previous proofs, where starting from the intializations, small learning rate ensures that certain noise alignments move towards 0 and others move away from 0 early in training. Hence, in this process, the signs do not change, and these sets always contain the same noises as $t$ progresses ∎
We now focus on the special case of the theorem and show that in expectation, our requirement on the generation noise prevents $\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{G,(t)},{\bm{w}}_{j}^{(t)})$ from getting too large.
Lemma C.1 implies the sets of noises to which we overfit or do not overfit remain unchanged during early training. Hence, we eliminate the possibility that noises move from one set to another, which simplifies the big picture. The technique is similar to what we did in Theorem 4.1, and we show that this also holds true for data augmented via generation.
For Theorem 4.2, the expectation is taken with respect to the underlying data (and noise) distributions. We first start with computing $\mathbb{E}\left[\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{G,(t)},{\bm{w}}_{j}^{(t)})\right]$. Expanding using the definition of NoiseAlign, we have that this quantity equals:
$\mathbb{E}\left[\frac{1}{|{\mathcal{I}}_{j,+}^{G,(t)}|}\left(\sum_{\begin{subarray}{c}i=1,\\ \bm{\xi}_{i}\in{\mathcal{I}}_{j,+}^{G,(t)}\end{subarray}}^{N}\left|\langle\nabla_{{\bm{w}}_{j}^{(t)}}{\mathcal{L}}({\bm{W}}^{(t)}),{\bm{\xi}}_{i}\rangle\right|+\sum_{\begin{subarray}{c}i=N+1,\\ \bm{\gamma}_{i}\in{\mathcal{I}}_{j,+}^{G,(t)}\end{subarray}}^{N_{\text{new}}}\left|\langle\nabla_{{\bm{w}}_{j}^{(t)}}{\mathcal{L}}({\bm{W}}^{(t)}),\bm{\gamma}_{i}\rangle\right|\right)\right]$
@@
where $p=\frac{N}{N_{new}}=\frac{1}{\alpha+k(1-\alpha)}$ measures the probability that the noise belongs to the original data instead of the synthetic data. The last equality is due to the fact that $\bm{\phi}_{i}\in{\mathcal{I}}_{j,+}^{G,(t)}$ are generated i.i.d and follow the same noise distribution $\mathcal{D}$} or ${\mathcal{D}}_{\bm{\gamma}}$.
Recall the following assumption:
$\mathbb{E}\left[l_{\bm{\gamma}}(t)\langle{\bm{w}}_{j}^{(t)},\bm{\gamma}\rangle^{2}\|\bm{\gamma}\|^{2}\right]$ $<(k+1)\mathbb{E}\left[l_{\bm{\xi}}(t)\langle{\bm{w}}_{j}^{(t)},\bm{\xi}\rangle^{2}\|\bm{\xi}\|^{2}\right]$
$\mathbb{E}\left[\frac{3}{N}l_{\bm{\gamma}}(t)\langle{\bm{w}}_{j}^{(t)},\bm{\gamma}\rangle^{2}\|\bm{\gamma}\|^{2}\right]$ $<(k+1)\mathbb{E}\left[\frac{3}{N}l_{\bm{\xi}}(t)\langle{\bm{w}}_{j}^{(t)},\bm{\xi}\rangle^{2}\|\bm{\xi}\|^{2}\right]$
$\mathbb{E}\left[\left|\langle\nabla_{{\bm{w}}_{j}^{(t)}}{\mathcal{L}}({\bm{W}}^{(t)}),\bm{\gamma}\rangle\right|\right]$ $<(k+1)\mathbb{E}\left[\left|\langle\nabla_{{\bm{w}}_{j}^{(t)}}{\mathcal{L}}({\bm{W}}^{(t)}),{\bm{\xi}}\rangle\right|\right],\quad\bm{\gamma}\sim{\mathcal{D}}_{\bm{\gamma}},\ \ {\bm{\xi}}\sim{\mathcal{D}}$
where the last inequality follows the expression of noise gradient. Again, note that ${\mathcal{I}}_{j,+}^{G,(t)}$ contains synthetic noises that follow the distribution ${\mathcal{D}}_{\bm{\gamma}}$ since the synthetic data points are generated independently. However, the noises in ${\mathcal{I}}_{j,+}^{U,(t)}$ are not independent due to upsampling and therefore do not follow the noise distribution in $\mathcal{D}$}.
To tackle this issue, we use the gradient alignment bound above and continue with Equation C
[…]
$=\$$\frac{1}{\alpha+k(1-\alpha)}\,\mathbb{E}\left[\left|\langle\nabla_{{\bm{w}}_{j}^{(t)}}{\mathcal{L}}({\bm{W}}^{(t)}),{\bm{\xi}}\rangle\right|\right]+\frac{(k^{2}-1)(1-\alpha)}{\alpha+k(1-\alpha)}\,\mathbb{E}\left[\left|\langle\nabla_{{\bm{w}}_{j}^{(t)}}{\mathcal{L}}({\bm{W}}^{(t)}),{\bm{\xi}}\rangle\right|\right]$
[…]
The last equality follows from the distribution of noises in the upsampled dataset: the slow-learnable $(1-\alpha)N$ data points are copied $k$ times, and from the gradient computation (first part of this theorem), each is learned $k$ times more, giving rise to the two $k$’s. The noises in each summation term follow $\mathcal{D}$}, as they are independent within the subset of summands.
Hence, we have shown the desired result:
$\mathbb{E}\left[\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{G,(t)},{\bm{w}}_{j}^{(t)})\right]<\mathbb{E}\left[\textbf{NoiseAlign}({\mathcal{I}}_{j,+}^{U,(t)},{\bm{w}}_{j}^{(t)})\right].$
At a high level, this suggests that starting at the weight ${\bm{w}}_{j}^{(t)}$, as long as the notion of small synthetic noise is satisfied, the expected gradient alignment with noises in the set where overfitting occurs is strictly lower when the dataset is augmented through generation. This reduced expected NoiseAlign leads to gradient updates that place less emphasis on fitting noise, thereby enabling the model to focus more effectively on learning the true features.
Consequently, although both upsampling and generation promote more uniform learning, their distinct interactions with noise critically influence generalization performance. This highlights the necessity of incorporating synthetic data and motivates our method.
@@
$B_{1}=B\frac{\alpha N}{\alpha N+k(1-\alpha)N}=B\frac{\alpha}{\alpha+k(1-\alpha)}.$ (12)
$B_{2}=B\frac{k(1-\alpha)N}{\alpha N+k(1-\alpha)N}=B\frac{k(1-\alpha)}{\alpha+k(1-\alpha)}.$ (13)
@@
$=\frac{\sigma^{2}}{B}\left(1+\frac{k(k-1)(1-\alpha)}{(\alpha+k(1-\alpha))^{2}}\frac{B}{N}\right)=I.$
@@ Appendix E Pseudocode
Algorithm 1 illustrates our method. We include Table 3 to compare it with other augmentation methods, including classical ones, naive diffusion-based generation, and mix-based methods such as mixup (Zhang et al. (2017)) and CutMix (Yun et al. (2019)).
Table 3: Comparison with other augmentation methods.
Feature Classical Aug (Flip, Crop, etc.) Mix-based (mixup/CutMix)
Core Objective Learn basic invariances Model regularization
Selection Strategy Applied randomly to all images Random pairing of images; augments all samples
Mechanism Geometric/color transformations Mix patches between real images
Computation Negligible: native GPU operations Low: simple arithmetic operations
Guarantee (Theory) No Yes
Primary Application Image classification; transfer learning Image classification; transfer learning
Key Advantage Fast, effective baseline Cheap and effective regularizer
Feature Diffusion-based Our Method
Core Objective Increase data quantity and diversity Reduce simplicity bias; promote balanced feature learning
Selection Strategy Selects samples not learned early in training
Mechanism Generates synthetic images and mixes them with real images Generates synthetic images close to real images
Computation High: generates multiple times for different conditional prompts Moderate: only generates a subset (30–40% of dataset)
Guarantee (Theory) No Yes
Primary Application Image classification; transfer learning Image classification; transfer learning
Key Advantage Can boost performance at a large compute cost Efficient: high gain from few samples; can be combined with prior methods
@@ F.1 Additional experimental settings
Datasets. The CIFAR10 dataset consists of 60,000 32 × 32 color images in 10 classes, with 6000 images per class. The CIFAR100 dataset is just like the CIFAR10, except it has 100 classes containing 600 images each. For both of these datasets, the training set has 50,000 images (5,000 per class for CIFAR10 and 500 per class for CIFAR100) with the test set having 10,000 images. Tiny-ImageNet comprises 100,000 images distributed across 200 classes of ImageNet Deng et al. (2009), with each class containing 500 images. These images have been resized to 64×64 dimensions and are in color. The dataset consists of 500 training images, 50 validation images, and 50 test images per class. For transfer learning experiments, we also used three fine-grained classification datasets. Flowers-102 (Nilsback & Zisserman, 2008) contains 8,189 images of flowers from 102 categories, with between 40 and 258 images per class. Aircraft (Maji et al., 2013) consists of 10,000 images across 100 aircraft model variants, with roughly 100 images per class. Stanford Cars (Krause et al., 2013) contains 16,185 images of 196 classes of cars, with classes defined at the level of make, model, and year.
Training on different datasets. From the setting in Andriushchenko & Flammarion (2022), we trained Pre-Activation ResNet18 on all datasets for 200 epochs with a batch size of 128. We used SGD with the momentum parameter 0.9 and set weight decay to 0.0005. We also fixed $\rho=0.1$ for SAM in all experiments unless explicitly stated. We used a linear learning rate schedule starting at 0.1. The learning rate is decayed by a factor of 10 after 50% and 75% epochs, i.e., we set the learning rate to 0.01 after 100 epochs and to 0.001 after 150 epochs. For transfer learning experiments, we fine-tuned a pre-trained ResNet18 on ImageNet-1K. On Flowers-102, we trained for 200 epochs with an initial learning rate of 0.001. On Aircraft and Stanford Cars, we trained for 100 epochs with an initial learning rate of 0.01.
Training with different architectures. We used the same training procedures for Pre-Activation ResNet18, VGG19, and DenseNet121. We directly used the official Pytorch Paszke et al. (2019) implementation for VGG19 and DenseNet121. For ViT Yuan et al. (2021), we adopted a Pytorch implementation at https://github.com/lucidrains/vit-pytorch. In particular, the hidden size, the depth, the number of attention heads, and the MLP size are set to 768, 8, 8, and 2304, respectively. We adjusted the patch size to 4 to fit the resolution of CIFAR10 and set both the initial learning rate and $\rho$ to 0.01. For both ConvNeXt-T and Swin-T, we used their official implementations but trained with SGD instead of Adam. The (learning rate, $\rho$) is set to (0.1, 0.05) for ConvNeXt-T and (0.01, 0.001) for Swin-T.
Hyperparameters. For Upsample Nguyen et al. (2024), we adopt their default hyper-parameters for separating epochs and set the upsampling factor $k$ to 2 as it yields the best performance. For our method, the upsampling factor $k$ is set to $5$ for CIFAR10 and CIFAR100 while that of TinyImageNet is set to $4$. Table 4 summarizes the upsampling factor along with the denoising steps at which the synthetic images are used to augment the base training set. For transfer learning experiments, we set $k$ to 2.
@@
Generating synthetic data with diffusion model. We use an open-source text-to-image diffusion model, GLIDE Nichol et al. (2021) as our baseline for image generation. We directly used the official Pytorch implementation at https://github.com/openai/glide-text2im. While generating, we set the time steps as 100 and the guidance scale as 3.0. For a k-way classification, we input the class names $C=\{c_{1},\dots,c_{k}\}$ with prompt $l=^{\prime}\text{a photo of}\{c_{i}\}^{\prime}$. For DiffuseMix, we used their official implementation with a single conditional prompt (“Autumn”).
@@
Figure 6: Examples of slow-learnable (row 1) and fast-learnable (row 3) samples in CIFAR-10. Rows 2 and 4 show the corresponding synthetic images generated with GLIDE (Section 4.4) using 50 denoising steps (FID = 10.67).